I need explanation and solution:
Rustam can paint a wall in 1 hour and Anvar can paint the same wall in 2 hours. How long would it take Rustam and Anvar to paint this wall if they work together?

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Rustam can paint in 1 hour
so he does 1 job in 1 hour
Anwar takes 2 hours
so he does 1/2 job in 1 hour
together they do 1+1/2 of the job in 1 hour
3/2 of the job in 1 hour
so for 1 job they together take 2/3 hours

Answer by richwmiller(17219)

x/1+x/2=1
2x/2+x/2=1
3x/2=1
x=2/3